# On Gauss-Kraïtchik formula for cyclotomic polynomials via symmetric functions

### v6

@@ 1. Introduction
for a constant $C>0$ explicitly given by a complicate expression. Bateman [2] proved that
@@
Gauss’ formula in Sections 356–357, p.p.633–638 of Disquitiones Arithmeticae (or p.p.425–428 of German translation [5]) states that if $d$ is an odd prime, then we can have
@@
where $\Psi_{d}(X)$ is a polynomial of degree (d1)/2(d-1)/2 and of integer coefficients beginning with two, $\Xi_{d}(X)$ is a polynomial of degree (d3)/2(d-3)/2 and of integer coefficients beginning with one, and $D=(-1)^{(d-1)/2}d$. Later, Kraïtchik proved that (6) holds for any odd squarefree $d$ with $\Psi_{d}(X)$ of degree $\varphi(d)/2$ and $\Xi_{d}(X)$ of degree $\varphi(d)/2-1$ in Tome I, p.p.124–129 of [9] (see also p.330 of [14]). Tome II, p.p.2–4 of [9] gives coefficients of $\Psi_{d}(X)$ and $\Xi_{d}(X)$ for $5\leq d\leq 101$ and Table 33 in p.p.445–451 of [14] extends this for $5\leq d\leq 149$. For example, we have
@@
Gauss-Kraïtchik formula is useful for studies of arithmetic properties of values of cyclotomic polynomials and values of the sum $\sigma(N)$ of divisors of $N$. For example, in [18] and [19], the author used Gauss’ formula to obtain upper bounds for odd perfect numbers of a certain form, where an upper bound for the ratio $\Xi_{d}(X)/\Psi_{d}(X)$ plays an important role.
@@
Let $d\geq 5$ be an odd squarefree integer, $0\leq n\leq d^{\prime}$, and $F=F_{d,n}=\max\{\varphi(f)/2:f\mid d,1<f\leq n\}\cup\{\left\lvert 1+\sqrt{D}\right\rvert/2\}$. Then,
@@
and, putting $G_{d,n}=\max\{\varphi(f)/2:f\mid d,1<f\leq n\}\cup\{\left\lvert 1+\sqrt{d}\right\rvert/2\}$,
@@
Let
(12) $U_{d}^{\pm}(X)=\prod_{1\leq k\leq d,(k/d)=\pm 1}(X-e^{2\pi ik/d}).$
Then we can write
@@
Moreover, we can easily see that $u_{d,n}=a_{d,n}+b_{d,n}\sqrt{D}$. Coefficients $u_{d,n}$ are what we consider actually. For example, if $d$ is an odd prime, then
@@
which can be confirmed using the Girard-Newton identity and the formula for quadratic Gauss sums, as we do later.
We also have an approximation for the ratio $\left\lvert\Xi_{d}(x)/\Psi_{d}(x)\right\rvert$, which generalizes Lemma 2.3 and the note in the corrigendum of [18].
@@
This theorem gives an error term of order $O(1/(x^{2}\sqrt{d}))$. We note that it would be possible to have an approximate formula with an error term of order $O(1/(x^{n}\sqrt{d}))$ for a positive integer $n$ if we have a formula of $u_{d,k}$ like (15) and (16) for each $k\leq n-1$. However, in order to obtain such formulae, we must process several cases according to values of quadratic residue symbols $(k/d)$ over $k=2,\ldots,n-1$ and the main term would be rather complicated.
Our argument depends on general relations of symmetric functions and evaluation of power sums of roots of unity. Coefficients $u_{d,n}$ of $U_{d}^{+}(X)$ can be represented as elementary symmetric polynomials. We can write such polynomials using power sums of roots of unity using the well-known Girard-Newton identity. The identity itself is rather complicated to obtain our upper bounds (but we still need it in order to determine signs of some quantities). Instead, we prove another specialized but simple identity. Combining with evaluation of power sums of roots of unity, we can bound sizes of coefficients $u_{d,n}$.
@@ 2. Power sums of roots of unity
If $\gcd(k,d)=1$, then $s_{d,k}=\sum_{(a/d)=1}\zeta_{d}^{ka}=\sum_{(a/d)=(k/d)}\zeta_{d}^{k}$ and hence $s_{d,k}=s_{d}^{(k/d)}=(\mu(d)+(k/d)\sqrt{D})/2$ as desired.
@@ 3. Elementary symmetric polynomials and power sums polynomials
We clearly have $\sum_{m=0}^{\infty}S^{(m)}T^{m}=-\prod_{i=1}^{\infty}(1+x_{i}T)$ and therefore

### v7

@@ 1. Introduction
for a constant $C>0$ explicitly given by a complicated expression. Bateman [2] proved that
@@
Gauss’ formula in Sections 356–357, pp. 633–638 of Disquitiones Arithmeticae (or pp. 436–440 of the translation [5]) states that if $d$ is an odd prime, then we can have
@@
where $\Psi_{d}(X)$ is a polynomial of degree (d1)/2(d-1)/2 and of integer coefficients beginning with two, $\Xi_{d}(X)$ is a polynomial of degree (d3)/2(d-3)/2 and of integer coefficients beginning with one, and $D=(-1)^{(d-1)/2}d$. Later, Kraïtchik proved that (6) holds for any odd squarefree $d$ with $\Psi_{d}(X)$ of degree $\varphi(d)/2$ and $\Xi_{d}(X)$ of degree $\varphi(d)/2-1$ in Tome I, pp. 124–129 of [9] (see also p.330 of [14]). Tome II, pp. 2–4 of [9] gives coefficients of $\Psi_{d}(X)$ and $\Xi_{d}(X)$ for $5\leq d\leq 101$ and Table 33 in pp. 445–451 of [14] extends this for $5\leq d\leq 149$. For example, we have
@@
Gauss-Kraïtchik formula gives representations of values of cyclotomic polynomials as norms in quadratic fields and is useful for studies of arithmetic properties of such values and the sum $\sigma(N)$ of divisors of $N$. For example, in [18] and [19], the author used Gauss’ formula to obtain upper bounds for odd perfect numbers of a certain form, where an upper bound for the ratio $\Xi_{d}(X)/\Psi_{d}(X)$ plays an important role.
@@
Let $d\geq 5$ be an odd squarefree integer, $0\leq n\leq d^{\prime}$, and $F=F_{d,n}=\max\left(\{\varphi(f)/2:f\mid d,1<f\leq n\}\cup\{\left\lvert 1+\sqrt{D}\right\rvert/2\}\right)$. Then,
@@
and, putting $G_{d,n}=\max\left(\{\varphi(f)/2:f\mid d,1<f\leq n\}\cup\{\left\lvert 1+\sqrt{d}\right\rvert/2\}\right)$,
@@
Our argument depends on general relations of symmetric functions and evaluation of power sums of roots of unity. Let
(12) $U_{d}^{\pm}(X)=\prod_{1\leq k\leq d,(k/d)=\pm 1}(X-e^{2\pi ik/d}),$
where $(\cdot/d)$ denotes the Jaoobi symbol. Then, for an odd squarefree integer $d$, we can write
@@
Moreover, we can easily see that $u_{d,n}=a_{d,n}+b_{d,n}\sqrt{D}$. These coefficients $u_{d,n}$ are the main objects of our study. For example, if $d$ is an odd prime, then
@@
which can be confirmed using the Girard-Newton identity and the formula for quadratic Gauss sums, as we do later. In general, coefficients $u_{d,n}$ of $U_{d}^{+}(X)$ can be represented as elementary symmetric polynomials. We can write such polynomials using power sums of roots of unity using the well-known Girard-Newton identity. The identity itself is rather complicated to obtain our upper bounds (but we still need it in order to determine signs of some quantities). Instead, we prove another specialized but simple identity. Combining with evaluation of power sums of roots of unity, we can bound sizes of coefficients $u_{d,n}$.
One of our motivation is quantification of how representations of values of cyclotomic polynomials as norms in quadratic fields constrain the relative sizes of conjugate embeddings. As we wrote before, the author used Gauss’ formula to obtain upper bounds in [18] and [19], for odd perfect numbers of a certain form, where an upper bound for the ratio $\Xi_{d}(X)/\Psi_{d}(X)$ plays an important role.
For example, We also have an approximation for the ratio $\left\lvert\Xi_{d}(x)/\Psi_{d}(x)\right\rvert$, which generalizes Lemma 2.3 and the note in the corrigendum of [18].
@@
This theorem gives an error term of order $O(1/(x^{2}\sqrt{d}))$ to clarify how representations of values of cyclotomic polynomials as norms in quadratic fields constrain the relative sizes of conjugate embeddings. We note that it would be possible to have an approximate formula with an error term of order $O(1/(x^{n}\sqrt{d}))$ for a positive integer $n$ if we have a formula of $u_{d,k}$ like (15) and (16) for each $k\leq n-1$, which would be useful to study other diophantine and arithmetic problems involving cyclotomic values or integers of special forms such as $(x^{m}-1)/(x-1)$ and $x^{m}-1$. However, in order to obtain such formulae, we must process several cases according to values of quadratic residue symbols $(k/d)$ over $k=2,\ldots,n-1$ and the main term would be rather complicated.
@@ 2. Power sums of roots of unity
If $\gcd(k,d)=1$, then $s_{d,k}=\sum_{(a/d)=1}\zeta_{d}^{ka}=\sum_{(a/d)=(k/d)}\zeta_{d}^{a}$ and hence $s_{d,k}=s_{d}^{(k/d)}=(\mu(d)+(k/d)\sqrt{D})/2$ as desired.
@@ 3. Elementary symmetric polynomials and power sums polynomials
We clearly have $\sum_{m=0}^{\infty}S^{(m)}T^{m}=\prod_{i=1}^{\infty}(1+x_{i}T)$ and therefore